Agenda/Minutes on 2020/10/16
CRY+VIS Interface Meeting
Participants: Takahashi, Sato, Hirata, Ikeda, Miyo, Hagiwara, Kimura, Aso, Tomaru, Ushiba, Akutsu
Cable bandling bellow BF.(Assy small hexagon for BF-CRYO.pdf)
- PF actuator
IP monitoring improvement in vacuum pumping -> auto-tracking system
- safe mode switching schedule of WAB
- Yano-san's schedule
Participants: Takahashi, Fabian, Hirata, Iguchi, Ishizaki, Miyo, Tanaka, Sato, Suguyama, Terrence, Aso
I wrote a document about the procedure to replace the LVDT cards after a failure: JGW-T2012128.
- It was checked by Alessandro Betolini from Nikhef and he says the method is correct. He provided minor comments that have been incorported to the new version of the document.
- He says they will review the design for ETpathfinder and that's our chance to give them feedback (i.e. Tanaka-san's findings and my concerns as a user).
- Knowledge transfer from Tanaka-san to Alessandro would be greatly beneficial for KAGRA. If you think it's suitable, would it be possible for someone to organize this, please?
- Tanaka-san will take some of our LVDT cards for hardware modification.
We need to survey those LVDT cards before he takes them away. When will we do this?
As far as I understand, given the old calibration factor, he has to give us a procedure to calculate a new calibration factor. Is this statement correct?
Before any other card breaks, we need to survey all of them. When will we do this?
Four hands are required for the measurements. For Type-B/Bp, Washimi-san and Fabian can do it.
After cleaning the PR area I have the impression we lost Hirose-san's LED lamps to check mirror cleanliness.
- We went to clean-up of each clean booth.
- I gave some feedback for the noise budget paper.
- We finished the cleaning with a lot of help.
- Hirata-san made M6 holes in some ballast masses.
- We removed some cables from chassis for the cleaning procedure.
- I put the fibre optic camera in the PR clean booth buffer area.
- We recovered the suspensions after blackout. They all went to ALIGNED state well.
- The temperature has decreased a lot inside the clean rooms, likely due to an increase of cold air circulation after turning all FFUs on. A couple of DC control loops in GAS filters are about to saturate. I've reported this to Miyoki-san and he turned on some heaters.
- After the blackout PR2 went to ALIGNED state well. PR3 and PRM oplevs were removed before recovery and I just checked they went to TWR_DAMPED state. I can turn the IM control system on by hand to check the health of IM-OSEMS.
- Just for the record for new comers: I realized PR use different amplitudes of reference signals for the LVDTs. According to labels attached to signals generators:
- PR2: 6 Vpp.
- PR3 and PRM: 7 Vpp.
- More two channels are not available in 1U chassis. 2U chassis are necessary with bridge cable.
- Cables for the primary coils can connected to the LVDT driver directly. Signals for secondary coils are exchanged inside the chassis.
- RSE Trial was finished on Oct. 13th.
- Preparation work is on going.
- Procedure documents.
- Criteria in the acceptance test.
The week from 10/19
- Takahashi: 10/19-23
The week from 10/26
- Hirata: 10/26-30
- On 2020/10/23(Fri)